OpenStudy (aravindg):

A 100 km/h wind blows nrmally against one wall of a house with an area of 50 m square.calculate force exerted on wall if the air moves parallel to wall after striking it and has a density of 1.134kg/m cube.

OpenStudy (anonymous):

The idea is that the momentum of the wind will be zero at the end (the wind will disperse in every direction so the vector sum is zero). m*v=F*t F=m*v/t =rho*V*v/t =rho*A*d*v/t =rho*A*(d/t)*v =rho*A*v^2 =(27.77777m/s)^2*50m^2*1.134kg/m^3 =43750 kg*m/s^2=43750 N. You can think that the mass flow per second is: (27.77777m/s)*50m^2*1.134kg/m^3=1575 kg/s and its "thrust" is mass flow times the velocity is (1575 kg/s)*27.7777m/s=43750 newtons.